OpenStudy (anonymous):

A trailer is rented at $30 per day plus a charge per mile of use. The trailer traveled 25 miles in one day and the total charge was $130. Which equation can be used to find the charge x, in dollars, per mile traveled? 30x + 25 = 130 30x + 25x = 130 30 + 25x = 130 30 + 25 = 130x

OpenStudy (anonymous):

30x + 25 = 130 30x + 25x = 130 30 + 25x = 130 30 + 25 = 130x

OpenStudy (anonymous):

the independent variable here are the miles

OpenStudy (anonymous):

we know the trailer cost $30 per day, nothing changes that what we don't know is the price per mile however, 15 miles and the $30 fee accounted for $130 we know it travelled 15 miles we don't know the cost for miles (x)

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

ok we have 15 miles and what will give us a fee?

OpenStudy (anonymous):

15 miles alone no fee. just miles

OpenStudy (anonymous):

Omg. I am not good at this. Math. Is. Terrible. I think you should just give up now...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

well the $30 fee does not depend on the mileage. if the trailer just sit there in the street and the sun shines on it. no milega and still $30

OpenStudy (anonymous):

so the 30 should NOT be multiplied with the "cost per mile"

OpenStudy (anonymous):

(cost per mile = x )

OpenStudy (anonymous):

the cost per mile should also not be multiplied with what you pay for everything. that would mean you treat the payment like miles and then compute cost per payment mile?? not the way to go! (so 130 also should not get x)

OpenStudy (anonymous):

well cost per mile times miles will give you the total mileage cost. if the cost per mile is $7, then 10 miles would cost - 10x costpermile, or 10x$7 = $70

OpenStudy (anonymous):

can you understand briefly, how the cost per mile is computed?

OpenStudy (anonymous):

Kind of..Sorry, I really am not good at this..:/

OpenStudy (anonymous):

your answer is this: 30 + 25x = 130 ^fixed cost for renting the trailer, this doesn't depend on miles so we don't multiply with miles 30 + 25x = 130 ^total payment, consisting of fixed cost and mileage cost 30 + 25x = 130 ^computation of cost due to mileage. more mileage = more cost. every mile cost x dollars. so 25*x gives cost for 25 miles 30 + 25x = 130